OpenStudy (anonymous):

F = -2px i - py j (p constant) on a particle. Show F is conservative and that p(x^2 + (y^2/2)) is a potential energy function for F. Find work done to move particle from (5,0) to (0,5)

OpenStudy (anonymous):

derivative of -2p x wrt to y is 0 derivative 0f -py wrt to x is 0 hey they are the same , hence conservative \[-2p \int x dx \] \[-p x^2 + g(y)\] \[-p \int y dy \] \[-p \frac{y^2}{2}+ h(x)\] \[p(x,y)=-p x^2 -p \frac{y^2}{2}\]

OpenStudy (anonymous):

p(x,y) at (0,5)-p(x,y) at (5,0) = work done

OpenStudy (anonymous):

Yup, 25p/2

OpenStudy (anonymous):

Now, if they weren't conservative , it would be a pain trying to remember stoke